Dear,

Here is a summary of CREAM-IV flight preparation status on the 21st of November at McMurdo.

  • Weather briefing shows no significant changes in the stratosphere. Still the weather pattern analysis shows summer weather pattern to be 2 days later than the last year.
  • All instruments were turned on except CAL HV from 8:30 AM for the End-to-End test in the Iridium connection mode.
  • Data were taken in a flight mode successfully during the test.
  • The reduced HSK data were received at SOC and displayed.

 

  • The commanding from the SOC was successful also.
  • Because of time, the parameter uploading through the Iridium couldn't be finished.  There was a delay in copying the parameter file from the cdps2 machine since the commanding syntax was not correct at the beginning. 
  • Will restart the uploading test from tomorrow morning 9 AM. But no connection from the SOC is necessary.
  • The Iridium test started 9 AM and ended at 11 AM.
  • The LPT with the CDM started and finished successfully.
  • After the LPT, playback of a set of yesterday's data done. Another set will have a playback tomorrow whenever possible.
  • The followings are to do list before the CPT in the next week.
  • Take muon runs for TCD/Chercam/CVD tuning if time available.
  • The top SCD data is unstable in the gain runs as we have seen before. Will be checked before the CPT with the CDM starts next Wednesday.
  • The calorimeter MB#7 temperature HSK display is incorrect as we have been seen. Will replace the board before the next CPT also.

Cheers, MooHyun.

UMD